Consider the following frequency table of observations on the random variable X.

Values 0 1 2 3 4 5
Observed Frequency 8 25 22 21 16 8

(a) Based on these 100 observations, is a Poisson distribution with a mean of 2.4 an appropriate model? Perform a goodness-of-fit procedure with α=0.05. Which of the following is the correct conclusion?

(b) Which of the following are the correct bounds on the P-value for this test.
Answer #1

Ho: distribution is poisson distribtuion

H1: distribution do not follow poisson distribution

total observation = ΣO =100      

expected proportion = e * µ^x / x!

expected frequency = expected proportion * total observation      
µ =    2.400

X observed (O) expected proportion = e-2.4* 2.4^x / x! expected frequency ( E ) (O-E)²/E
0 8 0.091 9.072 0.127
1 25 0.218 21.772 0.478
2 22 0.261 26.127 0.652
3 21 0.209 20.901 0.000
4 16 0.125 12.541 0.954
5 8 0.096 9.587 0.263

chi square test statistic,X² = Σ(O-E)²/E =   2.47
  
level of significance, α=   0.05              
Degree of freedom=k-2 =    6   -   1   =   5
Critical value =    11.0705   [ Excel function: =chisq.inv.rt(α,df) ]          

since, test stat < critical value, fail to reject Ho

there is not enough evidence that distribution follow poisson distribution..

b)

P value =   0.7804   [ excel function: =chisq.dist.rt(test-stat,df) ]      

0.05< p value <0.10